OpenStudy (anonymous):

Find the area of a circle whose circumference is 12pi ft. A. 113.04 square feet B. 24 square feet C. 144 square feet D. 108.12 square feet i think its C but not sure please help medal and fan

OpenStudy (tkhunny):

Circumference = \(12\pi\;ft\) Diameter = Circumference / \(\pi\) Radius = Diameter / 2 Area = \(\pi\cdot Radius^{2}\) Follow the path.

OpenStudy (anonymous):

I dont get it

OpenStudy (tkhunny):

Did you follow the path? 1) Calculate the Diameter.